At a prayer meeting, the disciples sing JAI-RAM JAI-RAM. The sound amplified by a loudspeaker comes back after reflection from a building at a distance of 80 m from the meeting. What maximum time interval can be kept between one JAI-RAM and the next JAI-RAM so that the echo does not disturb a listener sitting in the meeting. Speed of sound in air is 320 m s-1.

Grade:10

1 Answers

Deepak Patra
9 years ago
Sol. Here given S = 80 m × 2 = 160 m. v = 320 m/s So the maximum time interval will be t = 5/v = 160/320 = 0.5 seconds.